OpenStudy (anonymous):

As part of the technical crew for the circus, you need to set up the safety net for the human canon ball. He is launched at a 45 degree at a velocity of 25 m/s. Where will you place the net?

OpenStudy (johnweldon1993):

 So if he is launched at 25m/s at a 45 degree angle...we just need to find the horizontal component of that vector And we know that, in this case,the horizontal component an be found using the cosine function \[\large \cos(45) = \frac{x}{25m/s}\] So we know that \[\large x = 25m/s \times cos(45)\]

OpenStudy (anonymous):

Thanks! I got it now!!

OpenStudy (johnweldon1993):

No problem!